### v1

Abstract
For a subfield $\mathbb{K}$ of the field $\mathbb{C}$ of complex numbers, we consider curve and divisorial valuations on the algebra $\mathbb{K}[[x,y]]$ of formal power series in two variables with the coeficients in $\mathbb{K}$. We compute the semigroup Poincaré series and the classical Poincaré series of a valuation.
1 Introduction
A (discrete, rank one) valuation on the algebra ${\mathcal{O}}_{\mathbb{C}^{2},0}$ of germs of holomorphic functions in two variables is a map $\nu:{\mathcal{O}}_{\mathbb{C}^{2},0}\to\mathbb{Z}_{\geq 0}\cup\{+\infty\}$ such that
  1. 1)
    $\nu(0)=+\infty$;
  2. 2)
    $\nu(\lambda f)=\nu(f)$ for $f\in{\mathcal{O}}_{\mathbb{C}^{2},0}$, $\lambda\in\mathbb{C}^{*}:=\mathbb{C}\setminus\{0\}$;
  3. 3)
    $\nu(f_{1}+f_{2})\geq\min\{\nu(f_{1}),\nu(f_{2})\}$ for $f_{1},f_{2}\in{\mathcal{O}}_{\mathbb{C}^{2},0}$;
  4. 4)
    $\nu(f_{1}f_{2})=\nu(f_{1})+\nu(f_{2})$ for $f_{1},f_{2}\in{\mathcal{O}}_{\mathbb{C}^{2},0}$.
There are essentially two classes of valuations on ${\mathcal{O}}_{\mathbb{C}^{2},0}$: curve valuations and divisorial ones (see, e. g., [11]).
The notion of the Poincaré series of a collection $\{\nu_{i}\}$, $i=1,\ldots,r$, of valuations was introduced in [3]. For $r=1$ it reduces to
$P_{\nu}(t)=\sum_{v=0}^{\infty}\dim_{\mathbb{C}}(J^{\mathbb{C}}(v)/J^{\mathbb{C}}(v+1))\cdot t^{v}\in\mathbb{Z}[[t]]\;,$ (1)
where $J^{\mathbb{C}}(v):=\{f\in{\mathcal{O}}_{\mathbb{C}^{2},0}:\nu(f)\geq v\}$. The Poincaré series of one curve valuation was known to specialists for many years. Computations can be found, e. g., in [4]. The Poincaré series of one divisorial valuation was computed in [9]. The Poincaré series of a collection of curve valuations was computed in [5]. It was shown that this series coincides with the Alexander polynomial in several variables of the corresponding algebraic link. This implies that the Poincaré series determines the topology of the union of the corresponding curves.
In [7], there were considered valuations on the algebra ${\mathcal{E}}_{\mathbb{R}^{2},0}$ of germs of real analytic functions on $(\mathbb{R}^{2},0)$. (All valuations on ${\mathcal{E}}_{\mathbb{R}^{2},0}$ are restrictions of valuations on ${\mathcal{O}}_{\mathbb{C}^{2},0}$.) There were considered several versions of the Poincaré series (including the one defined in [3]). There was computed the Poincaré series for one curve valuation. In[8], there was computed the Poincaré series of one divisorial valuation in this setting.
2 Curve valuations on functions over a field
Let $\mathbb{K}$ be a subfield of the field $\mathbb{C}$ of complex numbers and let ${\mathcal{E}}_{\mathbb{K}^{2},0}$ be the algebra of germs of holomorphic functions in two variables with the Taylor coefficients from $\mathbb{K}$. A (discrete, rank one) valuation on ${\mathcal{E}}_{\mathbb{K}^{2},0}$ (defined as above) is the restriction of a valuation on ${\mathcal{O}}_{\mathbb{C}^{2},0}$: see, e. g., [10].
In the framework of the discussion below, without changes it is possible to consider, instead of the algebra ${\mathcal{E}}_{\mathbb{K}^{2},0}$, the algebra $\mathbb{K}[[x,y]]$ of formal power series in two variables. We shall refer to ${\mathcal{E}}_{\mathbb{K}^{2},0}$ because of some convenience of the language.
Let $(C,0)$ be a plane algebroid branch on $(\mathbb{C}^{2},0)$, that is $C$ (after an appropriate change of the coordinates in $\mathbb{C}^{2}$) is defined by
$x=x(\tau)=\tau^{m},\\ y=y(\tau)=\sum_{i\geq m}c_{i}\tau^{i}\in\mathbb{C}[[\tau]]\,.$ (2)
(We assume that the parametrization (2) is irreducible, i. e., the greatest common divisor of $m$ and of all $i$ such that $c_{i}\neq 0$ is equal to $1$.) The branch $C$ defines a valuation $\nu=\nu_{C}$ on the algebra ${\mathcal{O}}_{\mathbb{C}^{2},0}$ of holomorphic function germs in $x$ and $y$. For $f\in{\mathcal{O}}_{\mathbb{C}^{2},0}$, $\nu(f)$ is the degree of the leading term in the series $f(x(\tau),y(\tau))\in\mathbb{C}[[\tau]]$:
$f(x(\tau),y(\tau))=a(f)\cdot\tau^{\nu(f)}+\text{ terms of higher degree}\;,$ (3)
where $a(f)\neq 0$. If $f(x(\tau),y(\tau))\equiv 0$, $\nu(f)):=+\infty$. Let us consider its restriction to the subalgebra ${\mathcal{E}}_{\mathbb{K}^{2},0}\subset{\mathcal{O}}_{\mathbb{C}^{2},0}$. For $v\in\mathbb{Z}_{\geq 0}$, let $J^{\mathbb{K}}(v):=\{f\in{\mathcal{E}}_{\mathbb{K}^{2},0}:\nu(f)\geq v\}$. If all the quotients $J^{\mathbb{K}}(v)/J^{\mathbb{K}}(v+1)$ are finite-dimensional (over $\mathbb{K}$) the definition in [3] reduces to the following one:
Definition 1.
The $\mathbb{K}$-Poincaré series of the branch $C$ is
$P_{C}(t)(=P_{C}^{\mathbb{K}}(t))=\sum_{v=0}^{\infty}\dim_{\mathbb{K}}\left(J^{\mathbb{K}}(v)/J^{\mathbb{K}}(v+1)\right)\cdot t^{v}.$ (4)
Example 1.
If $\mathbb{K}=\mathbb{R}$, $P_{C}(t)$ is the classical Poincaré series of the valuation $\nu$ computed in [7].
Remark 1.
If $\mathbb{K}=\mathbb{C}$, all the coefficients of the series (4) are equal to 00 or $1$. This is not the case, in general, for a proper subfield $\mathbb{K}\subset\mathbb{C}$. In particular, if $\mathbb{K}=\mathbb{R}$ (the field of real numbers), all the coefficients are equal to 00, $1$, or $2$ ([7]).
Let $S^{\mathbb{K}}_{C}:=\{\nu(f):f\in{\mathcal{E}}_{\mathbb{K}^{2},0}\}\subset\mathbb{Z}_{\geq 0}$ be the semigroup of values of the valuation $\nu_{C}$ on ${\mathcal{E}}_{\mathbb{K}^{2},0}$.
Definition 2.
The $\mathbb{K}$-semigroup Poincaré series of the valuation $\nu$ is $P^{S}_{C}(t)=\sum_{v\in S^{\mathbb{K}}_{C}}t^{v}$.
Example 2.
For $\mathbb{K}=\mathbb{R}$ this is the semigroup Poincaré series considered (and computed) in [7].
Here we compute the $\mathbb{K}$-semigroup Poincaré series and the $\mathbb{K}$-Poincaré series for one curve or divisorial valuation.
3 The resolution process of a branch
Let $G$ be the Galois group of the extention $\mathbb{C}/\mathbb{K}$. i. e. the group of automorphisms of $\mathbb{C}$ which are trivial on $\mathbb{K}$. Let us consider the minimal $G$-invariant resolution process of the branch $C$ (by a sequence of blow-ups). This means the following. When we blow-up the origin in $\mathbb{C}^{2}$, we have got one of the following three possibilities.
  1. 1)
    There are finitely many different branches in the $G$-orbit of the strict transform $\widetilde{C}$ of the branch $C$, all of them are smooth and they intersect the exceptional divisor transversally at pairwise different points. We get (the minimal embedded $G$-invariant) resolution of $C$ and the process is finished.
  2. 2)
    The branches from the $G$-orbit of $\widetilde{C}$ (whose number may be infinite) intersect the exceptional divisor at finite number of points, but either they are not smooth, or do not intersect the exceptional divisor transversally, or (some) different branches from the orbit intersect it at the same point. In this case one has to continue the process.
  3. 3)
    The branches from the $G$-orbit of $\widetilde{C}$ intersect the exceptional divisor at infinitely many different points. In this case the process is terminated.
In Case 2 one has to blow-up (all) the intersection points of the branches from the $G$-orbit of $\widetilde{C}$ with the exceptional divisor (their number being finite). One gets one of the analogues the cases 1–3 again. Etc.
As the result we arrive to one of the following three situations.
  1. I
    There are finitely many branches in the $G$-orbit of $C$ and they are resolved at a certain step. The dual graph of the resolution looks like in Figure 1.
Figure 1: The resolution graph $\Gamma$ in case I (here $s=2$, $g=6$, see below).
The Galois group $G$ acts on the graph $\Gamma$ and the quotient $\check{\Gamma}$ by the action looks like in Figure 2 (and further on).
The graph $\Gamma$ contains a subgraph which geometrically coincides with the dual graph of a usual resolution of the curve $C$: the lower part of $\Gamma$ on Figure 1. (We do not consider this resolution below.) We use “short” notations (by greek letters) for verticies on this subgraph (and also for the corresponding verticies of the graph $\check{\Gamma}$). All other verticies are obtain from them by the action of elements of the Galois group $G$.
In Figure 1, $\sigma_{i}$, $i=0,\ldots,g$, (and also their images under the action of $G$) are the dead ends of the graph ($g$ is the number of the Puiseux pairs of the curve $C$), $\tau_{i}$, $i=1,\ldots,g$, (and also their images) are the rupture points of it, $\rho_{j}$, $j=1,\ldots,s$, ($s$ is equal to $\infty$ in Case II) are the splitting points corresponding to the changes of the isotropy subgroups for the $G$-action on $\Gamma$ (that is the splitting points for usual resolution of different branches from the orbit of $C$, there are finitely many of them in Case I infinitely many in Case II.
Figure 2 shows $\check{\Gamma}$ in the case when the process of the $\mathbb{K}$-resolution of $C$ is not finish when $C$ itself is resolved. Otherwise $\delta_{C}$ coincides with $\tau_{g}$.
Figure 2: The resolution graph $\check{\Gamma}$ in Case I (with arbitrary $g$ and $s$).
  1. II
    The described process goes on without termination. This means that we get an infinite sequence of blow-ups. The corresponding dual graph of the process looks like in Figure 3. One can say that $s$ is equal to infinity in this case.
    Figure 3: The resolution graph $\check{\Gamma}$ in Case II (with $s=\infty$).
  2. III
    The process is terminated at a certain step. This means that at this step the branches from the $G$-orbit of $\widetilde{C}$ intersect each of the new-born components of the exceptional divisor at infinitely many points. The dual graph of the process looks like in Figure 4.
    In the latter case, the valuation coincides with $n\nu$, where $\nu=\nu_{\rho_{s}}$ is the divisorial valuation defined by $E_{\rho_{s}}$ and $n$ is the intersection multiplicity $[E_{\rho_{s}},\widetilde{C}]$ of the strict transform $\widetilde{C}$ of the curve $C$ with the component $E_{\rho_{s}}$.
    Figure 4: The resolution graph $\check{\Gamma}$ in Case III.
Cases I–III take place in the following situations.
  1. I
    All the coefficients in the series (2) are from a finite extension of $\mathbb{K}$. This is equivalent to the condition that the orbit $GC$ has finitely many elements.
  2. II
    The coefficients all together are not from a finite extension of $\mathbb{K}$, but each of them is algebraic over $\mathbb{K}$.
  3. III
    There exists (at least) one coefficient transcendental over $\mathbb{K}$.
Remark 2.
Let $\mathbb{K}\subset\mathbb{C}$ (e. g., $\mathbb{K}=\mathbb{Q}$) and let $F\in\mathbb{K}[[x,y]]$ be irreducible. The local ring of the irreducible curve (over $\mathbb{K}$) defined by $F$ is $R=\mathbb{K}[[x,y]]/(F)$. The integral closure $\overline{R}$ in its quotient field is the valuation ring $R_{\omega}$ of a discrete valuation $\omega$, its residual field $\mathbb{K}^{\prime}$ is a finite algebraic extension of $\mathbb{K}$ and could be realized and therefore identified with a subfield of $\mathbb{C}$. The equation $F$ defines a reduced algebroid curve over the complex field $\mathbb{C}$, each branch corresponds to an irreducible factor of $F$ in $\mathbb{C}[[x,y]]$. Let $C$ be one of these branches. The valuation $\nu=\nu_{C}$ defined by $C$ on $\mathbb{K}[[x,y]]$ coincides with the valuation $\omega$ (seen as a valuation on $\mathbb{K}[[x,y]]$ forgeting the quotient by $F$) and the Poincaré series (over $\mathbb{K}$) defined by $\nu$ is the Poincaré series defined by $\omega$. Note that all the branches are equisingular, the minimal resolution of $\{F=0\}$ (in the complex setting) coincides with the $\mathbb{K}$-resolution of $\nu_{C}$ and all the valuations $\nu_{C}$, for complex branches $C$ of $\{F=0\}$, restrict to the same valuation $\omega$ in $\mathbb{K}[[x,y]]$. Moreover, if $G^{\prime}$ is the Galois group of the extension $\mathbb{K}\subset{\mathbb{L}}$, ${\mathbb{L}}$ the normal closure of $\mathbb{K}^{\prime}$ in $\mathbb{C}$, $G^{\prime}$ is the quotient of $G$ by the isotropy group of the field $\mathbb{L}$. This situation just corresponds to the case I described above.
4 Subgroups of the Galois group, subfields, divisors and curvettes
Further on we shall discuss only Cases I and II. Let $(E_{\sigma}\circ E_{\delta})$, $\sigma,\delta\in\Gamma$, be the intersection matrix of the components $E_{\sigma}$ (infinite one in Case II). The diagonal entry $E_{\sigma}\circ E_{\sigma}$ is the self-intersection number of the component $E_{\sigma}$ on the final modification in Case I and on a modification at a step, after which the points of $E_{\sigma}$ are not blown-up, in Case II. One has $E_{\sigma}\circ E_{\sigma}<0$. The non-diagonal entry $E_{\sigma}\circ E_{\delta}$ is equal to 1 (respectively to 00) if these components intersect (do not intersect respectively).
Let $(m_{\sigma,\delta}):=-(E_{\sigma}\circ E_{\delta})^{-1}$. In Case II one can compute the entry $m_{\sigma,\delta}$ at a sufficiently late step (after which it stabilizes). In other words one can compute the intersection matrix and the inverse of it at each step and use their stabilization.
The number $m_{\sigma,\delta}$ can be interpreted as the intersection number of some germs of curves on $(\mathbb{C}^{2},0)$, so-called curvettes. A curvette correponding to the component $E_{\sigma}$ is the blow-down $C_{\sigma}$ of a complex-analytic smooth curve $\gamma_{\sigma}$ (in Case II on a modification where $E_{\sigma}$ stabilizes) transversal to the component $E_{\sigma}$ at a smooth point of the exceptional divisor, that is not an intersection point with other components. One can show that $m_{\sigma,\delta}$ is equal to the intersection number $C_{\sigma}\circ C_{\delta}$.
For $\sigma\in\Gamma$, let $m_{\sigma}$ be the value of $\nu_{C}$ on the left-hand side of an equation $f_{\sigma}=0$ defining the curvette $C_{\sigma}$, i. e., the intersection number $C\circ C_{\sigma}$. One can see that, in Case I, $m_{\sigma}$ is equal to $m_{\sigma,\delta_{C}}$ for $\delta_{C}$ such that, on the surface of the resolution, the strict transform of the branch $C$ intersects the component $E_{\delta_{C}}$ and, in Case II, $m_{\sigma}$ is equal to $m_{\sigma,\delta}$ for $\delta$ on the subgraph of the usual resolution of the curve $C$ far enough from the initial component of the modification.
Isotropy subgroups and subfields.
Let $G_{\sigma}$ be the isotropy subgroup of the component $E_{\sigma}$, i. e., the isotropy subgroup of the vertex $\sigma$ in $\Gamma$. The subgroup $G_{\sigma}$ has a finite index in the Galois group $G$. Let $\mathbb{K}_{\sigma}\subset\mathbb{C}$ be the invariant subfield of the group $G_{\sigma}$. One can see that, for $g\in G$, $G_{g\sigma}=gG_{\sigma}g^{-1}$, $\mathbb{K}_{g\sigma}=g\mathbb{K}_{\sigma}$. For all the vertices $\sigma\in\Gamma$ inbetween the initial one $\sigma_{0}=$1 and $\rho_{1}$ ($\rho_{1}$ included) the subgroups $G_{\sigma}$ and the subfields $\mathbb{K}_{\sigma}$ are the same: $G_{\sigma}=G=G_{0}$, $\mathbb{K}_{\sigma}=\mathbb{K}=\mathbb{K}_{0}$. For all vertices $\sigma\in\Gamma$ inbetween the vertex $\rho_{j}$ and $\rho_{j+1}$ ($\rho_{j}$ excluded, $\rho_{j+1}$ included) the isotropy subgroups $G_{\sigma}$ and the corresponding subfields $\mathbb{K}_{\sigma}$ are the same: $G_{\sigma}=G_{j}$, $\mathbb{K}_{\sigma}=\mathbb{K}_{j}$. In Case I, for all the vertices $\sigma\in\Gamma$ betwin $\rho_{s}$ and $\delta_{C}$ one has $G_{\sigma}=G_{s}$, $\mathbb{K}_{\sigma}=\mathbb{K}_{s}$. For vertices from $G$-orbits of the discussed ones, the corresponding subgroups are conjugate to $G_{\sigma}$ and the subfields are obtained from $\mathbb{K}_{\sigma}$ by the shift by an element of the Galois group $G$. One has $G=G_{0}\supset G_{1}\supset\ldots\supset G_{s}$, $\mathbb{K}=\mathbb{K}_{0}\subset\mathbb{K}_{1}\subset\ldots\subset\mathbb{K}_{s}$. One has $[K_{j}:K_{j-1}]=[G_{j-1}:G_{j}]$.
Standard charts.
For computations, we shall use coordinate systems compatible with the actions of the Galois groups. This means the following. There are two standard charts covering the initial component of the resolution process: with the coordinates $u=x,w=y/x$ or $u=y,w=x/y$ respectively. If, on a previously created component one makes the blow-up at the point $(u,w)=(0,a)$ (in a standard chart), one has two standard charts covering the new-born component: with the coordinates $u^{\prime}=u,w^{\prime}=(w-a)/u$ and $u^{\prime}=w-a,w^{\prime}=u/(w-a)$ respectively.
On all the components $E_{\sigma}$ of the exceptional divisor except $E_{\rho_{j}}$ and those from its G-orbit, on each step the blow-up is made at one point with the coordinates $(0,a)$ in a standard chart with $a\in\mathbb{K}_{\sigma}$. In this case it can be convenient to change the coordinate $w$ by $\widetilde{w}=w/(w-a)$ or $\widetilde{w}=(w-a)/w$ so that the coordinates of the intersection points with the components on the geodesic between $\sigma_{0}$ and $\delta_{C}$ become equal to 00 and $\infty$. On the component $E_{\rho_{j}}$, the blow-ups are made at a certain point $(0,a)$ with $a\in\mathbb{K}_{j}\setminus\mathbb{K}_{j-1}$ (that is not from $\mathbb{K}_{\rho_{j}}=\mathbb{K}_{j-1}$) and at the points from its $G_{j-1}$-orbit. In this case a change as above is inappropriate.
Special curvettes.
In what follows, we shall use curvettes of special type. In a standard chart covering an affine part of $E_{\sigma}$, let the (smooth and transversal) curve $\gamma_{\sigma}$ be defined over the field $\mathbb{K}_{\sigma}$. In other words, the curve $\gamma_{\sigma}$ is invariant under the action of $G_{\sigma}$. In particular one can take the curve $\{u=\tau$, $w=c=\text{const}\}$ in a standart chart with $c\in\mathbb{K}_{\sigma}$ ($u=0$ is the equation of $E_{\sigma}$ in the chart).
Definition 3.
A $\mathbb{K}_{\sigma}$-curvette at the component $E_{\sigma}$ is the blow-down $C_{\sigma}$ of a curve $\gamma_{\sigma}$ described above.
Remark 3.
One can see that $C_{\sigma}$ has a parametrization like (2) with all the coefficients from $\mathbb{K}_{\sigma}$. A $\mathbb{K}_{\sigma}$-curvette at the component $E_{\sigma}$ can be defined by an equation $f_{\sigma}=0$ with $f_{\sigma}\in{\mathcal{E}}_{\mathbb{K}_{\sigma}^{2},0}$.
Definition 4.
A $G_{\sigma}$-curvette at the component $E_{\sigma}$ is the union of the curves $gC_{\sigma}$ ($C_{\sigma}$ is a $\mathbb{K}_{\sigma}$-curvette at $E_{\sigma}$) over representatives of the $G_{\sigma}$-classes in $G$.
One can show that a $G_{\sigma}$-curvette at the component $E_{\sigma}$ can be defined by an equation $F_{\sigma}=0$ with $F_{\sigma}\in{\mathcal{E}}_{\mathbb{K}^{2},0}$. In fact one can take $F_{\sigma}=\prod_{[g]\in G/G_{\sigma}}(gf_{\sigma})$, where the product is over representatives of the $G_{\sigma}$-classes in $G$.
Remark 4.
Somewhat similar type of symmetries of the resolution graph was considered in [6]. In that case one had an action of a group on the graph, but not on the underlying space $\mathbb{C}^{2}$ or on the space of functions. The series computed in this paper and therein are different. In fact the setting studied here was one of the motivations for [6].
5 The semigroup Poincaré series of a curve valuation
Let $\nu(=\nu_{C})$ be the valuation on ${\mathcal{E}}_{\mathbb{K}^{2},0}$ defined by a curve $C\subset(\mathbb{C}^{2},0)$ and assume that $\nu_{C}$ is a valuation of the type I or II described in Section 3. Let $P^{S}_{C}(t)=\sum_{m\in S^{\mathbb{K}}_{\nu}}t^{m}$ be the semigroup Poincaré series of $\nu_{C}$ (see Definition 2).
For $\sigma\in\check{\Gamma}$, let $M_{\sigma}$ be the value of $\nu$ on the left-hand side of the equation $F_{\sigma}=0$ definig a $G_{\sigma}$-curvette at the component $E_{\sigma}$. One can see that $M_{\sigma}=\sum_{[g]\in G/G_{\sigma}}m_{g\sigma}$. (Compare with the definition of $M_{\sigma}$ in [7], where $\mathbb{K}=\mathbb{R}$.)
The aim of this section is to prove the following statement.
Theorem 1.
$P^{S}_{C}(t)=\frac{\prod_{i=1}^{g}(1-t^{M_{\tau_{i}}})}{\prod_{i=0}^{g}(1-t^{M_{\sigma_{i}}})}\;.$
Let $S_{C}$ be the usual semigroup of the complex curve $C$, i. e., $S_{C}=\{\nu(f):f\in{\mathcal{O}}_{\mathbb{C}^{2},0}\}$. The set of multiplicities $\{m_{\sigma_{0}},\ldots,m_{\sigma_{g}}\}$ is the minimal set of generators of $S_{C}$. Let $e_{i}:=\gcd\{m_{\sigma_{0}},\,\ldots,m_{\sigma_{i}}\}$ for $i=0,1,\ldots,g$ and let $N_{i}=e_{i-1}/e_{i}$ for $i=1,2,\ldots,g$. (One has $m_{\tau_{i}}=N_{i}m_{\sigma_{i}}$ for $i=1,2,\ldots,g$: see, e. g., [12] or [2].)
For $\sigma\in\Gamma$, let $\pi_{\sigma}:(X_{\sigma},D_{\sigma})\to(\mathbb{C}^{2},0)$ be the minimal modification of $(\mathbb{C}^{2},0)$ such that $E_{\sigma}\subset D_{\sigma}$. In particular, $E_{\sigma}$ is the last exceptional component appearing in $X_{\sigma}$ and is produced by blowing-up at a point $p_{\sigma}$ of a previous component. For $f\in{\mathcal{O}}_{\mathbb{C}^{2},0}$, we will denote by $e_{\sigma}(f)$ the multiplicity of the strict transform by $\pi_{\sigma}$ of the curve $\{f=0\}$ at the point $p_{\sigma}$. Notice that $e_{\sigma}(f)$ coincides with the intersection multiplicity of the strict transform of $\{f=0\}$ and $E_{\sigma}$ in the surface $X_{\sigma}$.
For $j=1,\ldots,s$, let $\ell_{j}:=[\mathbb{K}_{j}:\mathbb{K}_{j-1}]=[G_{j-1}:G_{j}]:=|\frac{G_{j-1}}{G_{j}}|$.
Lemma 1.
Let $\omega\in\check{\Gamma}$ be such that $\rho_{d}<\omega\leq\rho_{d+1}$ and $\tau_{p}\leq\omega<\tau_{{p}+1}$. One has
$M_{\omega}=\sum_{i=0}^{{p}}k_{i}M_{\sigma_{i}}+\sum_{j=1}^{d}r_{j}(\ell_{j}-1)M_{\rho_{j}}$
for some non negative integers $k_{i}$ $(0\leq i\leq{p})$ and positive integers $r_{j}$ $(1\leq j\leq d)$.
Proof..
Let us assume that $\tau_{q-1}<\rho_{d}\leq\tau_{q}$ (see Figure 5 below). For $i=0,\ldots,g$, let $M_{i}:=M_{\sigma_{i}}$, $m_{i}:=m_{\sigma_{i}}$ and let $\varphi_{i}=0$ be the equation of a curvette at the component $E_{\sigma_{i}}$.
Let $\omega$ be such that $\rho_{d}<\omega\leq\rho_{d+1}$. Then one has
$M_{\omega}$ $=m_{\omega}+([G_{0}:G_{d}]-[G_{1}:G_{d}])e_{\rho_{1}}(\varphi_{\omega})m_{\rho_{1}}+([G_{1}:G_{d}]-[G_{2}:G_{d}])e_{\rho_{2}}(\varphi_{\omega})m_{\rho_{2}}+$
$\qquad\qquad+\cdots+([G_{d-1}:G_{d}]-1)e_{\rho_{d}}(\varphi_{\omega})m_{\rho_{d}}=$
$=m_{\omega}+\sum_{j=1}^{d}([G_{j-1}:G_{d}]-[G_{j}:G_{d}])e_{\rho_{j}}(\varphi_{\omega})m_{\rho_{j}}=m_{\omega}+\sum_{j=1}^{d}L^{d}_{j}e_{\rho_{j}}(\varphi_{\omega})m_{\rho_{j}}$
where $L^{d}_{j}=[G_{j-1}:G_{d}]-[G_{j}:G_{d}]$, $1\leq j\leq d$. In the same way, for $i=q,\ldots,{p}$ one has
$M_{i}=m_{i}+\sum_{j=1}^{d}L^{d}_{j}e_{\rho_{j}}(\varphi_{i})m_{\rho_{j}}\,.$ (5)
Figure 5: The position of the vertices.
Moreover, it is known that $m_{\omega}\in\langle m_{0},\,\ldots,m_{p}\rangle$ and so $m_{\omega}=k_{0}m_{0}+\sum_{i=1}^{p}k_{i}m_{i}$, for some $k_{0}\geq 0$ and $0\leq k_{i}<N_{i}$ for $i\geq 1$ (which are unique with this restrictions). Then one can write
$M_{\omega}=\sum_{0}^{p}k_{i}m_{i}+\sum_{j=1}^{d}L^{d}_{j}e_{\rho_{j}}(\varphi_{\omega})m_{\rho_{j}}$ (6)
Let us assume that $q<{p}$. For $j\leq d$ one has $e_{\rho_{j}}(\varphi_{\omega})/e_{\tau_{{p}-1}}(\varphi_{\omega})=e_{\rho_{j}}(\varphi_{{p}})/e_{\tau_{{p}-1}}(\varphi_{p})$, moreover $e_{\tau_{{p}-1}}(\varphi_{p})=1$, so $e_{\rho_{j}}(\varphi_{\omega})=e_{\tau_{{p}-1}}(\varphi_{\omega})e_{\rho_{j}}(\varphi_{p})$. It is known that $e_{i}=e_{\tau_{i}}(\varphi_{C})$, $i\leq g$ and $\varphi_{C}$ is an equation of the complex curve $C$. As a consequence, for $i\leq p$, one has $N_{i}=e_{i-1}/e_{i}=e_{\tau_{i-1}}(\varphi_{\omega})/e_{\tau_{i}}(\varphi_{\omega})$ and so
$e_{\rho_{j}}(\varphi_{\omega})=N_{p}e_{\tau_{p}}(\varphi_{\omega})e_{\rho_{j}}(\varphi_{p})\;.$
The same computations applied to $\varphi_{i}$ (instead of $\varphi_{\omega}$) and $\varphi_{i-1}$ (for $\tau_{i}>\tau_{i-1}>\rho_{j}$) show that $e_{\rho_{j}}(\varphi_{i})=N_{i-1}e_{\tau_{i-1}}(\varphi_{i})e_{\rho_{j}}(\varphi_{i-1})=N_{i-1}e_{\rho_{j}}(\varphi_{i-1})$.
Using this equality in the rigth part of the equation (6) and also the expression of $M_{i}$ above, one gets:
$\sum_{j=1}^{d}L^{d}_{j}e_{\rho_{j}}(\varphi_{\omega})m_{\rho_{j}}$ $=N_{p}e_{\tau_{p}}(\varphi_{\omega})\sum_{j=1}^{d}L^{d}_{j}e_{\rho_{j}}(\varphi_{p})m_{\rho_{j}}=$
$=k_{p}\sum_{j=1}^{d}L^{d}_{j}e_{\rho_{j}}(\varphi_{p})m_{\rho_{j}}+k^{\prime}_{p}\sum_{j=1}^{d}L^{d}_{j}e_{\rho_{j}}(\varphi_{p})m_{\rho_{j}}$
with $k^{\prime}_{p}=N_{p}e_{\tau_{p}}(\varphi_{\omega})-k_{p}\geq 1$, and so
$M_{\omega}=\sum_{0}^{{p}-1}k_{i}m_{i}+k_{p}M_{p}+k^{\prime}_{p}\sum_{j=1}^{d}L^{d}_{j}e_{\rho_{j}}(\varphi_{p})m_{\rho_{j}}\;.$
In the same way, using repeatedly that $e_{\rho_{i}}(\varphi_{j})=N_{i-1}e_{\rho_{j}}(\varphi_{i-1})$ for $i\geq q$ one gets:
$M_{\omega}=(\sum_{0}^{q-1}k_{i}m_{i})+\sum_{i=q}^{p}k_{i}M_{i}+r_{d}\sum_{j=1}^{d}L^{d}_{j}e_{\rho_{j}}(\varphi_{q})m_{\rho_{j}}$ (7)
for some integer $r_{d}\geq 1$. Now, using that $L^{d}_{j}=[G_{d-1}:G_{d}]L^{d-1}_{j}$ for $j\leq d-1$ in the last summand of the above equality (7), one has
$r_{d}\sum_{j=1}^{d}L^{d}_{j}e_{\rho_{j}}(\varphi_{q})m_{\rho_{j}}=r_{d}[G_{d-1}:G_{d}]\sum_{j=1}^{d-1}L^{d-1}_{j}e_{\rho_{j}}(\varphi_{q})m_{\rho_{j}}+r_{d}L^{d}_{d}e_{\rho_{d}}(\varphi_{q})m_{\rho_{d}}\;.$
Using that $L^{d}_{d}=[G_{d-1}:G_{d}]-1$, $e_{\rho_{j}}(\varphi_{q})=e_{\rho_{j}}(\varphi_{\rho_{d}})$, $e_{\rho_{d}}(\varphi_{\rho_{d}})=1$ and the corresponding expression for $M_{\rho_{d}}$ in the same terms as at the beginning of the proof for $M_{\omega}$, the last equality can be written as:
$r_{d}([G_{d-1}:G_{d}]-1)\left(\sum_{j=1}^{d-1}L^{d-1}_{j}e_{\rho_{j}}(\varphi_{\rho_{d}})m_{\rho_{j}}+m_{\rho_{d}}\right)+r_{d}\sum_{j=1}^{d-1}L^{d-1}_{j}e_{\rho_{j}}(\varphi_{q})m_{\rho_{j}}=$
$=r_{d}([G_{d-1}:G_{d}]-1)M_{\rho_{d}}+r_{d}\sum_{j=1}^{d-1}L^{d-1}_{j}e_{\rho_{j}}(\varphi_{q})m_{\rho_{j}}\;.$
Thus, taking into account that $\ell_{j}=[G_{j-1}:G_{j}]$ and the last equality, the equation (7) for $M_{\omega}$ becomes:
$M_{\omega}=\sum_{i=q}^{p}k_{i}M_{i}+r_{d}(\ell_{d}-1)M_{\rho_{d}}+\left(\sum_{0}^{q-1}k_{i}m_{i}+r_{d}\sum_{j=1}^{d-1}L^{d-1}_{j}e_{\rho_{j}}(\varphi_{q})m_{\rho_{j}}\right)\;.$ (8)
If $d=1$, one has $m_{i}=M_{i}$ for $i=0,\ldots,q-1$ and the proof is finished. Otherwise, we apply the induction hypothesis to the last two summands in (8) (in brackets). Notice that the factor $r_{d}\geq 1$ does not affects the proof above and the terms $e_{\rho_{j}}(\varphi_{q})$ (corresponding to a curvette at $E_{\sigma_{q}}$) could be treated (arithmetically) as we did with $e_{\rho_{j}}(\varphi_{\omega})$ for $j<d$. As a consequence, we have:
$M_{\omega}$ $=\sum_{i:\rho_{1}<\tau_{i}\leq\tau_{p}}k_{i}M_{i}+\sum_{i:\tau_{i}\leq\rho_{1}}k_{i}m_{i}+\sum_{j=1}^{d}r_{j}(\ell_{j}-1)M_{\rho_{j}}$
$=\sum_{i=0}^{{p}}k_{i}M_{i}+\sum_{j=1}^{d}r_{j}(\ell_{j}-1)M_{\rho_{j}}$
for some integers $r_{j}\geq 1$ ($1\leq j\leq d$). ∎
Remark 5.
For $\sigma\in\check{\Gamma}$, $\tau_{k}\leq\sigma<\tau_{k+1}$ it is known that $m_{\sigma}\in\langle m_{\sigma_{0}},\ldots,m_{\sigma_{k}}\rangle$ or equivalently that $e_{k}$ divides $m_{\sigma}$ (see [12]. Taken into account the expression for $M_{i}$ given in equation (5), this implies that $e_{i-1}$ divides $\sum_{j=1}^{d}L^{d}_{j}e_{\rho_{j}}(\varphi_{i})m_{\rho_{i}}$ and so $\gcd(M_{0},\,\ldots,M_{i})=\gcd(m_{0},\,\ldots,m_{i})=e_{i}$ and
$N_{i}=e_{i-1}/e_{i}=\gcd(M_{0},\,\ldots,M_{i-1})/\gcd(M_{0},\,\ldots,M_{i})\,.$
Lemma 2.
For $i=1,\ldots,g$ one has $(N_{i}-1)M_{\sigma_{i}}\notin\langle M_{\sigma_{0}},\ldots,M_{\sigma_{i-1}}\rangle$, $N_{i}M_{\sigma_{i}}\in\langle M_{\sigma_{0}},\ldots,M_{\sigma_{i-1}}\rangle$ and $N_{i-1}M_{\sigma_{i-1}}<M_{\sigma_{i}}$. In particular, $\{M_{\sigma_{0}},\ldots,M_{\sigma_{g}}\}$ is the minimal set of generators of $S=\langle M_{\sigma_{0}},\ldots,M_{\sigma_{g}}\rangle$ and, moreover, $S$ is the semigroup of values of a complex analytic irreducible curve.
Proof..
Note that, as a consequence of the Remark 5, $(N_{i}-1)M_{i}\notin\langle M_{0},\ldots,M_{i-1}\rangle$.
Let us prove that for $i=1,\ldots,g$, one has that $N_{i-1}M_{i-1}<M_{i}$. If $\rho_{d}\leq\tau_{i-1}<\tau_{i}$, the equation (5) for $M_{i}$ together with the equality $e_{\rho_{i}}=N_{i-1}e_{\rho_{i}}(\varphi_{i-1})$ for $i\leq d$ implie that
$M_{i}$ $=m_{i}+\sum_{j=1}^{d}L^{d}_{j}e_{\rho_{j}}(\varphi_{i})m_{\rho_{j}}=m_{i}+N_{i-1}\sum_{j=1}^{d}L^{d}_{j}e_{\rho_{j}}(\varphi_{i-1})m_{\rho_{j}}>$
$>N_{i-1}\left(m_{i-1}+\sum_{i=j}^{d}L^{d}_{j}e_{\rho_{j}}(\varphi_{i-1})m_{\rho_{j}}\right)=N_{i-1}M_{i-1}\;.$
If $\tau_{i-1}<\rho_{i}\leq\tau_{i}$ the equation (5) for $M_{i-1}$ is $M_{i-1}=m_{i-1}+\sum_{j=1}^{d^{\prime}}L^{d^{\prime}}_{j}e_{\rho_{j}}(\varphi_{i-1})m_{\rho_{j}}$ for some $d^{\prime}<d$. Thus, using the equality
$L^{d}_{j}=[G_{d-1}:G_{d}]L^{d-1}_{j}=\ell_{d-1}L^{d-1}_{j}=\ell_{d-1}\cdots\ell_{d^{\prime}}L^{d^{\prime}}_{j}$
one gets
$M_{i}$ $=m_{i}+\sum_{j=1}^{d}L^{d}_{j}e_{\rho_{j}}(\varphi_{i})m_{\rho_{j}}>N_{i-1}\left(m_{i-1}+\sum_{j=1}^{d}L^{d}_{j}e_{\rho_{j}}(\varphi_{i-1})m_{\rho_{j}}\right)$
$>N_{i-1}\left(m_{i-1}+\sum_{j=1}^{d^{\prime}}L^{d^{\prime}}_{j}e_{\rho_{j}}(\varphi_{i-1})m_{\rho_{j}}\right)=N_{i-1}M_{i-1}\;.$
It is known that the proved properties implies that the semigroup $S=\langle M_{\sigma_{0}},\ldots,M_{\sigma_{g}}\rangle$ is the semigroup of a complex irreducible plane curve (see [1]).
Proposition 1.
The set of integers $\{M_{\sigma_{0}},\,\ldots,M_{\sigma_{g}}\}$ is the minimal set of generators of the semigroup of values $S^{\mathbb{K}}_{C}$ of the valuation $\nu$ on ${\mathcal{E}}_{\mathbb{K}^{2},0}$.
Proof..
By Lemma 2 it is enough to prove that $S^{\mathbb{K}}_{\nu}$ is generated by $M_{\sigma_{0}},\,\ldots,M_{\sigma_{g}}$. It is clear that the semigroup $S^{\mathbb{K}}_{\nu}$ is generated by the integers $M_{\sigma}$, for all $\sigma\in\check{\Gamma}$. Thus one has to show that $M_{\sigma}\in\langle M_{\sigma_{0}},\ldots,M_{\sigma_{g}}\rangle$ for any $\sigma$.
Let $[\alpha,\beta]$ denote the geodesic in $\check{\Gamma}$ between the vertices $\alpha$ and $\beta$. If $\omega\in[\sigma_{0},\delta_{C}]$, then Lemma 1 gives
$M_{\omega}=\sum_{i=0}^{h}k_{i}M_{\sigma_{i}}+\sum_{j=1}^{d}r_{j}(\ell_{j}-1)M_{\rho_{j}}$
where $\rho_{d}<\omega\leq\rho_{d+1}$, $\tau_{h}\leq\omega<\tau_{h+1}$ and $r_{j},k_{i}$ are non-negative integers. The same equation applied to $\rho_{j}$ gives us that $M_{\rho_{j}}\in\langle M_{0},\,\ldots,M_{i}\rangle$ if $\tau_{i}\leq\rho_{j}<\tau_{i+1}$. In particular, $M_{\rho_{j}}\in\langle M_{0},\,\ldots,M_{h}\rangle\subset\langle M_{0},\,\ldots,M_{g}\rangle$
If $\omega\in[\sigma_{i},\tau_{i}]$ for some $i=1,\ldots,g$, then $\nu(\varphi_{\omega})=km_{\sigma_{i}}$ for a positive integer $k$. For the conjugate $g\omega$, $g\in G$, one has that $g{\omega}\in[g{\sigma}_{i},g{\tau}_{i}]$ and so $\nu(g{\varphi_{\omega}})=\nu_{g{C}}(\varphi_{\omega})=km_{g{\sigma}_{i}}$. Thus, in this case one has that $M_{\omega}=\nu(\sum g\varphi_{\omega})=k(\sum m_{g\omega})=kM_{\sigma_{i}}$. Note that $M_{\tau_{i}}=N_{i}M_{\sigma_{i}}$.
Proof..
(Of Theorem 1) It follows directly from Proposition 1 and Lemma 2 taking into account that $M_{\tau_{i}}=N_{i}M_{\sigma_{i}}$ for all $i\geq 1$. ∎
Remark 6.
As we have proved, $S^{\mathbb{K}}_{\nu}$ is the semigroup of a complex analytic branch $C^{\prime}$ (and so $P^{S}_{C}(t)$ is the usual Poincaré serie of it) different from $C$ if $m_{\rho_{1}}<m_{\tau_{g}}$. We do not know a geometric reason for the existence of the curve $C^{\prime}$.
6 The classical Poincaré series
Theorem 2.
In Cases I and II, one has
$P_{C}(t)$ $=$ $P_{C}^{S}(t)\cdot\prod_{j=1}^{s}\frac{1-t^{\ell_{j}M_{\rho_{j}}}}{1-t^{M_{\rho_{j}}}}$
$=$ $\frac{\prod_{i=0}^{g}(1-t^{M_{\tau_{i}}})}{\prod_{i=0}^{g}(1-t^{M_{\sigma_{i}}})}\cdot\prod_{j=1}^{s}\frac{1-t^{\ell_{j}M_{\rho_{j}}}}{1-t^{M_{\rho_{j}}}}$
($s=\infty$ in Case II).
Proof..
Let $P_{C}(t)=\sum_{v=0}^{\infty}a_{v}t^{v}$. For a collection $\{k_{\sigma}\}$, $\sigma\in\check{\Gamma}$, $k_{\sigma}\in\mathbb{Z}_{\geq 0}$, let ${\mathcal{E}}_{\mathbb{K}^{2},0}^{\{k_{\sigma}\}}$ be the set of germs $f\in{\mathcal{E}}_{\mathbb{K}^{2},0}$ such that the intersection multiplicity of the strict transform of the zero-level set $\{f=0\}$ with the component $E_{\sigma}$ of the exceptional divisor $D$ is equal to $k_{\sigma}$ and, moreover, this strict transform intersects $D$ only at smooth points of the total transform $\pi^{-1}(C)$ of the curve $C$. (Pay attention that the intersection multiplicity of the strict transform of $\{f=0\}$ with the component $E_{g\sigma}$, $g\in G$, is also equal to $k_{\sigma}$.) Let $\nu(\{k_{\sigma}\}):=\sum_{\sigma\in\check{\Gamma}}k_{\sigma}M_{\sigma}$. One can see that $\nu(\{k_{\sigma}\})=\nu(f)$ for any $f\in{\mathcal{E}}_{\mathbb{K}^{2},0}^{\{k_{\sigma}\}}$.
Remark 7.
Without loss of generality, we shall use the assumption that, for a fixed $v=\nu(\{k_{\sigma}\})$, the latter condition holds automatically. This can be achieved if one makes sufficiently many additional blow-ups at intersection points of the components of the total transform $\pi^{-1}(C)$.
Let $F^{\{k_{\sigma}\}}$ be the image of ${\mathcal{E}}_{\mathbb{K}^{2},0}^{\{k_{\sigma}\}}$ in the quotient $J^{\mathbb{C}}(v)/J^{\mathbb{C}}(v+1)\cong\mathbb{C}$ with $v=\nu(\{k_{\sigma}\})$. Insted of writing that $f\in{\mathcal{E}}_{\mathbb{K}^{2},0}^{\{k_{\sigma}\}}$ (and explaining which $\{k_{\sigma}\}$ is considered), we shall sometimes write that $f$ corresponds to the representation $v=\sum_{\sigma\in\check{\Gamma}}k_{\sigma}M_{\sigma}$ or simply to the expression $\sum_{\sigma\in\check{\Gamma}}k_{\sigma}M_{\sigma}$.
One has:
  1. 1)
    $\bigcup_{\{k_{\sigma}\}:\nu(\{k_{\sigma}\})=v}F^{\{k_{\sigma}\}}=J^{\mathbb{K}}(v)/J^{\mathbb{K}}(v+1);$
  2. 2)
    for each collection $\{k_{\sigma}\}$, $F^{\{k_{\sigma}\}}$ is the complement to a (finite) arrangement of (proper) vector $\mathbb{K}$-subspaces in the vector $\mathbb{K}$-space $\overline{F}^{\{k_{\sigma}\}}$ (the linear span of $F^{\{k_{\sigma}\}}$).
Let $d^{\{k_{\sigma}\}}$ be the dimension (over $\mathbb{K}$) of the vector space $\overline{F}^{\{k_{\sigma}\}}$. One obviously has
$a_{v}=\max_{\{k_{\sigma}\}:\nu(\{k_{\sigma}\})=v}d^{\{k_{\sigma}\}}.$
For $j=1,2,\ldots,s+1$ ($s$ may be equal to +$\infty$+ ∞), let $A_{j}$ be the set of collections $\{k_{\sigma}\}$ such that $k_{\sigma}=0$ for all $\sigma$ on the geodesic $[\rho_{j},\delta_{C}]$ from $\rho_{j}$ to $\delta_{C}$ (including the ends). For $j=s+1$ (if $s<+\infty$), we assume this geodesic to be empty. Let the series $P^{(j)}_{C}(t)=\sum_{v=0}^{\infty}a^{(j)}_{v}t^{v}$ be defined by
$a^{(j)}_{v}=\max_{\{k_{\delta}\}\in A_{j}:\nu(\{k_{\sigma}\})=v}d^{\{k_{\sigma}\}}.$
One obviously has $P_{C}(t)=P^{(s+1)}_{C}(t)$ if $s<+\infty$ and, if $s=+\infty$ (i. e., if the process of the $\mathbb{K}$-resolution of the curve $C$ is infinite), $P_{C}(t)$ is, in a natural sense, the limit of $P^{(j)}_{C}(t)$ for $j\to+\infty$.
For a function $f\in{\mathcal{E}}_{\mathbb{K}^{2},0}$, let $\widetilde{f}=f\circ\pi$ be the lifting of $f$ to the surface of the resolution. In order to compute (or evaluate) the dimension $d^{\{k_{\sigma}\}}$, we shall regularly use the following construction. Let $h\in{\mathcal{E}}_{\mathbb{K}^{2},0}$ be a fixed function with $\nu(h)=\nu(\{k_{\sigma}\})$. (In fact we shall take a fixed function from ${\mathcal{E}}_{\mathbb{K}^{2},0}^{\{k_{\sigma}\}}$.) For each function $f\in{\mathcal{E}}_{\mathbb{K}^{2},0}^{\{k_{\sigma}\}}$ we shall consider the ratio $\Psi={\widetilde{f}}/{\widetilde{h}}$. The function $\Psi$ is regular and different from zero at the intersection point of the exceptional divisor with the strict transform of the curve $C$. The set of values of these ratios for all $f\in{\mathcal{E}}_{\mathbb{K}^{2},0}^{\{k_{\sigma}\}}$ is isomorphic to the set $F^{\{k_{\sigma}\}}$. This follows from the following arguments. The set $F^{\{k_{\sigma}\}}$ is just the set of the coefficients $a(f)$ in Equation (3) for $f\in{\mathcal{E}}_{\mathbb{K}^{2},0}^{\{k_{\sigma}\}}$. The coefficient $a(f)$ is equal to the coefficient $a(h)$ (a fixed number) mutiplied by the value of the ratio $\Psi$.
If $\rho_{1}\neq\sigma_{0}$, the statement follows from the following six lemmas (and two Propositions). The necessary changes in the case $\rho_{1}=\sigma_{0}$ will be discussed later.
Lemma 3.
$a^{(1)}_{v}\neq 0$ if and only if $v\in S^{\mathbb{K}}_{C}$.
Proof..
Obviously $a^{(1)}_{v}=0$ for $v\not\in S^{\mathbb{K}}_{C}$. Any $v\in S^{\mathbb{K}}_{C}$ can be represented in the form $v=\sum_{i=0}^{g}k_{i}M_{\sigma_{i}}$ (Proposition 1). Any function $f$ corresponding to this representation gives a non-zero element in the corresponding $F^{\{k_{\sigma}\}}$. Therefore $d^{\{k_{\sigma}\}}\geq 1$ and $a^{(1)}_{v}\geq 1$. ∎
Lemma 4.
For each $v$, $a^{(1)}_{v}\leq 1=[\mathbb{K}_{0}:\mathbb{K}_{0}]$. As a consequence, $a^{(1)}_{v}=1=[\mathbb{K}_{0}:\mathbb{K}_{0}]$ if and only if $v\in S^{\mathbb{K}}_{C}$.
Proof..
Let $\{k_{\sigma}\}\in A_{1}$ and let $h$ be a fixed function from ${\mathcal{E}}_{\mathbb{K}^{2},0}^{\{k_{\sigma}\}}$. For $f\in{\mathcal{E}}_{\mathbb{K}^{2},0}^{\{k_{\sigma}\}}$, the ratio $\Psi={\widetilde{f}}/{\widetilde{h}}$ has no zeroes or poles on all the components from the union of the geodesics between $E_{\rho_{1}}$ and $E_{g\delta_{C}}$ for all $g\in G=G_{0}$. Therefore this is a (non-zero) constant on them. This constant is invariant with respect to $G$ and therefore belongs to $\mathbb{K}$. ∎
Lemmas 3 and 4 obviously imply the following statement.
Proposition 2.
One has
$P^{(1)}_{C}(t)=P^{S}_{C}(t)=\frac{\prod_{i=1}^{g}(1-t^{M_{\tau_{i}}})}{\prod_{i=0}^{g}(1-t^{M_{\sigma_{i}}})}\;.$
Lemma 5.
For each $j\geq 1$ and any $v$, $a^{(j)}_{v}\leq[\mathbb{K}_{j-1}:\mathbb{K}_{0}]$.
Proof..
The proof is essentially the same as the one of Lemma 4. For $\{k_{\sigma}\}\in A_{j}$ and a fixed function $h\in{\mathcal{E}}_{\mathbb{K}^{2},0}^{\{k_{\sigma}\}}$, the ratio $\Psi={\widetilde{f}}/{\widetilde{h}}$ ($f\in{\mathcal{E}}_{\mathbb{K}^{2},0}^{\{k_{\sigma}\}}$) has no zeroes or poles on all the components from the union of the geodesics between $E_{\rho_{j}}$ and $E_{g\delta_{C}}$ for all $g\in G_{j-1}$. Therefore this is a (non-zero) constant on them. This constant is invariant with respect to $G_{j-1}$ and therefore belongs to $\mathbb{K}_{j-1}$. ∎
Lemma 6.
Let $v\in S^{\mathbb{K}}_{C}$ be of the form $v=\ell M_{\rho_{j}}+b$ with $b\in S^{\mathbb{K}}_{C}$ and the maximal possible $\ell$ in this representation, assume that $\ell\leq\ell_{i}-1$. We also suppose that $v$ is not of the form $M_{\varkappa}+b^{\prime}$ with $\varkappa>\rho_{j}$ and $b^{\prime}\in S^{\mathbb{K}}_{C}$. (The corresponding case will be anayzed in Lemma 8.) Then $a_{v}^{(j+1)}=\ell[\mathbb{K}_{j-1}:\mathbb{K}_{0}]+a_{b}^{(j)}$.
Proof..
We shall use the coordinate $w$ on $E_{\rho_{j}}$ (from the corresponding standard coordinate chart) such that $w=\infty$ at the intersection point with the previous component on the the geodesic between $\sigma_{0}$ and $\rho_{j}$. Let $w=z_{0}\in\mathbb{K}_{j}\setminus\mathbb{K}_{j-1}$ be a point on $E_{\rho_{j}}$ blown-up in the process of the resolution of the curve $C$. One has $\mathbb{K}_{j}=\mathbb{K}_{j-1}[z_{0}]$ and $1$, $z_{0}$, $z_{0}^{2}$, …, $z_{0}^{\ell_{j}-1}$ is a basis of the vector space $\mathbb{K}_{j}$ over $\mathbb{K}_{j-1}$.
Let us represent a collection $\{k_{\sigma}\}$ with $\nu(\{k_{\sigma}\})=v$ in the form $k_{\sigma}=k^{1}_{\sigma}+k^{2}_{\sigma}$, where $k^{1}_{\rho_{j}}=\ell^{\prime}$ ($\ell^{\prime}\leq\ell$), $k^{1}_{\sigma}=0$ for $\sigma\neq\rho_{j}$, $k^{2}_{\rho_{j}}=0$. A function $f\in{\mathcal{E}}_{\mathbb{K}^{2},0}^{\{k_{\sigma}\}}$ can be represented as $f=f_{1}f_{2}$ with $f_{r}\in{\mathcal{E}}_{\mathbb{K}^{2},0}^{\{k^{r}_{\sigma}\}}$, $r=1,2$. Let $h_{r}$, $r=1,2$, be fixed functions from ${\mathcal{E}}_{\mathbb{K}^{2},0}^{\{k^{r}_{\sigma}\}}$. For certainty, as $h_{1}$ we shall take the $\ell^{\prime}$-th power of the left hand side of the equation of the $G_{j-1}$-curvette at the point $w=1$ of the component $E_{\rho_{j}}$. Let $h=h_{1}h_{2}$. One has $\Psi={\widetilde{f}}/{\widetilde{h}}=\Psi_{1}\Psi_{2}$, where $\Psi_{r}={\widetilde{f}_{r}}/{\widetilde{h}_{r}}$. The function $\Psi_{2}$ is constant on the union of all the components $E_{\sigma}$ for $\sigma$ from the geodesics between $\rho_{j}$ and $g\delta_{C}$ with $g\in G_{j-1}$ (including both end points of the geodesic). Its value on them belongs to $\mathbb{K}_{j-1}$ and the maximal possible dimension (over $\mathbb{K}$) of the set of its values is $a_{b}^{(j)}$. The function $\Psi_{1}$ is constant on the union of all the components $E_{\sigma}$ for $\sigma$ from the geodesic between $\rho_{j}$ and $g\delta_{C}$ with $g\in G_{j}$ excluding $\rho_{j}$ itself. Its restriction to the component $E_{\rho_{j}}$ is equal to
$\Psi_{1}|_{E_{\rho_{j}}}=\frac{c\left(\prod\limits_{p=1}^{\ell^{\prime}}(w-w_{p})\right)}{(w-1)^{\ell^{\prime}}}\,,$
where $w_{1},\ldots,w_{\ell^{\prime}}$ are the intersection points of the strict transform of the curve $\{f_{1}=0\}$ with the component $E_{\rho_{j}}$. All these intersection points $w_{p}$ belong to $\mathbb{K}_{j-1}$. This follows from the fact that the set of intersection points of the strict transform of the curve $\{f_{1}=0\}$ with $E_{\rho_{j}}$ is $G_{j}$-invariant, whereas the orbit of a point not belonging to $\mathbb{K}_{j-1}$ consists of $\ell_{j}$ points. Moreover, its value at $w=\infty$, that is $c$, belongs to $\mathbb{K}$. (This follows from the fact that this ratio is constant on the union of the components from the geodesic $[\sigma_{0},\rho_{j})$.) Thus up to the costant $(z_{0}-1)^{\ell^{\prime}}$ (which does not influence the dimensions) the value of $\Psi$ on the union of the components from the geodesic $(\rho_{j},\delta_{C}]$ is constant and is equal to $c\prod_{p=1}^{\ell^{\prime}}(z_{0}-w_{p})=c(z_{0}^{\ell^{\prime}}-(w_{1}+\ldots+w_{\ell^{\prime}})z_{0}^{\ell^{\prime}-1}+\ldots\pm(w_{1}\cdot\ldots\cdot w_{\ell^{\prime}}))$. Therefore the value of the ratio $\Psi$ on this union is equal to $c(\alpha_{\ell^{\prime}}z_{0}^{\ell^{\prime}}+\alpha_{\ell^{\prime}-1}z_{0}^{\ell^{\prime}-1}+\alpha_{1}z_{0}+\alpha_{0})$, where $\alpha_{0}$, …, $\alpha_{\ell^{\prime}-1}$ are generic points of $\mathbb{K}_{j-1}$ (because we can choose arbitrarily $w_{1},\,\ldots,w_{\ell^{\prime}}$ in $\mathbb{K}_{j-1}$) and $\alpha_{\ell^{\prime}}$ (also a point of $\mathbb{K}_{j-1}$) is the value of the ratio $\Psi_{2}$. The (maximal) dimension of the set of values of $\alpha_{\ell^{\prime}}$ is $a_{b^{*}}^{(j)}$, where $b^{*}=v-\ell^{\prime}M_{\rho_{j}}$. Since $1$, $z_{0}$, …$z_{0}^{\ell^{\prime}}$ are linear independent over $\mathbb{K}_{j-1}$, the total dimension is equal to $\ell^{\prime}[\mathbb{K}_{j-1}:\mathbb{K}_{0}]+a_{b^{*}}^{(j)}$. Thus it is maximal only for $\ell^{\prime}=\ell$ and is equal to $\ell[\mathbb{K}_{j-1}:\mathbb{K}_{0}]+a^{(j)}_{b}$.
Lemma 7.
If $\rho_{j}<\varkappa\leq\delta_{C}$ (that is $\varkappa$ is on the geodesic between $\rho_{j}$ and $\delta_{C}$ and $\varkappa\neq\rho_{j}$), then $a_{M_{\varkappa}}^{(j+1)}=[\mathbb{K}_{j}:\mathbb{K}_{0}]$.
Proof..
Any collection $\{k_{\sigma}\}$ with $\nu(\{k_{\sigma}\})=M_{\varkappa}$ obviously belongs to $A_{j+1}$. According to Lemma 5, $a_{M_{\varkappa}}^{(j+1)}\leq[\mathbb{K}_{j}:\mathbb{K}_{0}]$. Let us take the collection $\{k^{0}_{\sigma}\}$ with $k^{0}_{\varkappa}=1$ and $k^{0}_{\sigma}=0$ for $\sigma\neq\varkappa$. One has $\nu(\{k^{0}_{\sigma}\})=M_{\varkappa}$. Let $f=0$ be an equation of a $G_{\varkappa}(=G_{j})$-curvette intersecting the component $E_{\varkappa}$ at the point with the coordinate $\alpha\in\mathbb{K}_{j}$. Here we assume that the coordinate is equal to zero at the intersection point with the component on the geodesic to $\delta_{C}$ and to $\infty$ at the one with the component on the geodesic to $\sigma_{0}$. Let $g$ be one of these functions, say, with $\alpha=1$. The ratio $\Psi={\widetilde{f}}/{\widetilde{g}}$, for $f\in{\mathcal{E}}^{\{k^{0}_{\sigma}\}}_{\mathbb{K}^{2},0}$, has no zeroes or poles on all the components of each of the indicated geodesics and thus is constant on each of them. On the component $E_{\varkappa}$ this ratio is of the form $c\frac{w-\alpha}{w-1}$ with $c\in\mathbb{K}$ and (arbitrary!) $\alpha\in\mathbb{K}_{j}$. The value on $\delta_{C}$ (equal to the value at $w=0$, i.e. $c\alpha$) is an arbitrary point of $\mathbb{K}_{j}$. Therefore $d^{\{k^{0}_{\sigma}\}}=[\mathbb{K}_{j}:\mathbb{K}_{0}]$ and $a_{M_{\varkappa}}^{(j+1)}\geq[\mathbb{K}_{j}:\mathbb{K}_{0}]$. ∎
Lemma 8.
Let $\varkappa$ be as in Lemma 7. Due to Lemma 1, one has that $M_{\varkappa}=(\ell_{j}-1)M_{\rho_{j}}+b$ with $b\in S^{\mathbb{K}}_{C}$. Then $a_{b}^{(j)}=[\mathbb{K}_{j-1}:\mathbb{K}_{0}]$.
Proof..
Due to Lemma 1, one has $M_{\varkappa}=b^{\prime}+\sum_{q=1}^{j}(\ell_{q}-1)M_{\rho_{q}}$ with $b^{\prime}\in S^{\mathbb{K}}_{C}$. Let $c_{k}:=b^{\prime}+\sum_{q=1}^{k}(\ell_{q}-1)M_{\rho_{q}}$, $k=1,\ldots,j$. We have to show that $a^{(j)}_{c_{j-1}}=[\mathbb{K}_{j-1}:\mathbb{K}_{0}]$. We shall show that $a^{(k)}_{c_{k-1}}=[\mathbb{K}_{k-1}:\mathbb{K}_{0}]$ for all $k$. We shall use induction in $k$. For $k=1$ this is trivial: $a^{(1)}_{c_{0}}=1=[\mathbb{K}_{0}:\mathbb{K}_{0}]$. Due to Lemma 6, we have $a^{(k)}_{c_{k-1}}=(\ell_{k-1}-1)[\mathbb{K}_{k-2}:\mathbb{K}_{0}]+a^{(k-1)}_{c_{k-2}}$. By the induction assumption $a^{(k-1)}_{c_{k-2}}=[\mathbb{K}_{k-2}:\mathbb{K}_{0}]$ and therefore $a^{(k)}_{c_{k-1}}=\ell_{k-1}[\mathbb{K}_{k-2}:\mathbb{K}_{0}]=[\mathbb{K}_{k-1}:\mathbb{K}_{0}]$. ∎
Lemmas 58 imply the following statement.
Proposition 3.
One has
$P^{(j+1)}(t)$ $=$ $P^{(j)}(t)\cdot(1+t^{M_{\rho_{j}}}+t^{2M_{\rho_{j}}}+\ldots+t^{(\ell_{j}-1)M_{\rho_{j}}})$
$=$ $P^{(j)}(t)\cdot\frac{1-t^{\ell_{j}M_{\rho_{j}}}}{1-t^{M_{\rho_{j}}}}.$
Proof..
Equation (3) is equivalent to
$a_{v}^{(j+1)}=a_{v}^{(j)}+a_{v-M_{\rho_{j}}}^{(j)}+a_{v-2M_{\rho_{j}}}^{(j)}+\ldots+a_{v-(\ell_{j}-1)M_{\rho_{j}}}^{(j)}\,.$ (10)
If $v\in S^{\mathbb{K}}_{C}$ cannot be represented as $M_{\rho_{j}}+b$ with $b\in S^{\mathbb{K}}_{C}$, one has $a_{v}^{j+1}=a_{v}^{j}$, what is compatible with (10). Let $v=\ell M_{\rho_{j}}+b$ with $b\in S^{\mathbb{K}}_{C}$, $11\leq\ell<\ell_{j}-$ 1, and $v$ cannot be represented as $(\ell+1)M_{\rho_{j}}+b^{\prime}$ with $b^{\prime}\in S^{\mathbb{K}}_{C}$. In this case Equation (10) reduces to
$a_{v}^{(j+1)}=a_{v}^{(j)}+a_{v-M_{\rho_{j}}}^{(j)}+a_{v-2M_{\rho_{j}}}^{(j)}+\ldots+a_{v-(\ell-1)M_{\rho_{j}}}^{(j)}+a_{b}^{(j)}\,$ (11)
(all other summands are equal to zero). Since, for $10\leq\ell^{\prime}\leq\ell-$ 1, $v-\ell^{\prime}M_{\rho_{j}}=M_{\rho_{j}}+b^{*}$ with $b^{*}\in S^{\mathbb{K}}_{C}$ and, due to Lemma 7, $a_{M_{\rho_{j}}}^{(j)}=[\mathbb{K}_{j-1}:\mathbb{K}_{0}]$, all the summands in (11) except the last one are equal to $[\mathbb{K}_{j-1}:\mathbb{K}_{0}]$. Thus in this case (10) follows from Lemma 6.
Let $v=(\ell_{j}-1)M_{\rho_{j}}+b$ with $b\in S^{\mathbb{K}}_{C}$. There are two possibilities.
  1. 1)
    The value $v$ can be represented as $M_{\varkappa}+b^{\prime}$ with $b^{\prime}\in S^{\mathbb{K}}_{C}$, $\varkappa>\rho_{j}$. In this case (10) follows from Lemmas 7 and  8.
  2. 2)
    The value $v$ cannot be represented in the indicated form. In this case (10) follows from Lemma 6 again.
Remark 8.
One can see that the proof of Proposition 3 gives a little bit more strong statement which we shall use in Section 7. Let $A^{\prime}_{j}$ be the set of collections $\{k_{\sigma}\}$ such that $k_{\sigma}=0$ for all $\sigma$ on the geodesic from $\rho_{j}$ to $\delta_{C}$ excluding (!) $\rho_{j}$ and including $\delta_{C}$. For $j=s+1$ (if $s<+\infty$), we assume this geodesic to be empty. Let the series $\widetilde{P}^{(j)}_{C}(t)=\sum_{v=0}^{\infty}\widetilde{a}^{(j)}_{v}t^{v}$ be defined by
$\widetilde{a}^{(j)}_{v}=\max_{\{k_{\delta}\}\in A^{\prime}_{j}:\nu(\{k_{\sigma}\})=v}d^{\{k_{\sigma}\}}.$
Then one has
$\widetilde{P}^{(j)}_{C}(t)=P^{(j+1)}(t)=P^{(j)}(t)\cdot\frac{1-t^{\ell_{j}M_{\rho_{j}}}}{1-t^{M_{\rho_{j}}}}.$ (12)
Propositions 2 and 3 imply the statement of Theorem 2 in the case under consideration: $\rho_{1}\neq\sigma_{0}$. In the case $\rho_{1}=\sigma_{0}$, Lemmas 58 and thus Proposition 3 hold for $j\geq 2$. Proposition 2 does not hold. However, it is possible to compute the series $P^{(2)}_{C}(t)$ directly using Lemmas 58. Thus the statement of Theorem 2 holds in this case as well. ∎
Remark 9.
Any series $\zeta(t)=1+c_{1}t+c_{2}t^{2}+\cdots\in\mathbb{Z}[[t]]$ can be in a unique way represented as $\zeta(t)=\prod_{m=1}^{\infty}(1-t^{m})^{s_{m}}$ with $s_{m}\in\mathbb{Z}$. We say that the series is cyclotomic if this product/ratio of binomials is finite. There is a natural generalization of this notion for series in sevearl variables. Ussually in previous computations (say e.g. [5]) the Poincaré series appeared to be cyclotomic. This is not the case for the classical Poincaré series $P_{C}(t)$ in case II, i.e. when $s=\infty$.
Remark 10.
Let us assume that the curve valuation is of type I. Let $c$ be the conductor of the semigroup $S^{\mathbb{K}}_{C}$ (i.e the minimal element of $S^{\mathbb{K}}_{C}$ such that $c+\mathbb{Z}_{\geq 0}\subset S^{\mathbb{K}}_{C}$). It is known that $c$ can be expresed as
$c=\sum_{i=1}^{g}(N_{i}-1)M_{\sigma_{i}}-M_{\sigma_{0}}+1\;.$
Let $\Delta:=c+\sum_{j=1}^{s}(\ell_{j}-1)M_{\rho_{j}}$ and $\ell=\ell_{1}\cdots\ell_{s}=[\mathbb{K}_{s}:\mathbb{K}]$. The computation of the Poincaré series gives that $a_{v}=\dim J^{\mathbb{K}}(v)/J^{\mathbb{K}}(v+1)=\ell$ for all $v\geq\Delta$ and $a_{\Delta-1}<\ell$. Moreover, one has the following symmetry property for all $v\in\mathbb{Z}$:
$\dim\left({J^{\mathbb{K}}(v)}\left/{J^{\mathbb{K}}(v+1)}\right.\right)+\dim\left({J^{\mathbb{K}}(\Delta-1-v)}\left/{J^{\mathbb{K}}(\Delta-v)}\right.\right)=\ell\;$ (13)
Following the notations of Remark 2, the integral closure $\overline{R}$ of the local ring $R=\mathbb{K}[[x,y]]/(F)$ of the algebroid curve defined by $F$ is a valuation ring and is isomophic to $\mathbb{K}^{\prime}[[u]]$ ($u$ is an uniformizing parameter). One has:
Corollary 1.
Let
$\Delta=\sum_{i=1}^{g}(N_{i}-1)M_{\sigma_{i}}-M_{\sigma_{0}}+1+\sum_{j=1}^{s}(\ell_{j}-1)M_{\rho_{j}}\;.$
Then, the conductor ideal ${\mathcal{C}}(\overline{R}/R)=\{z\in\overline{R}:z\overline{R}\subset R\}$ is the ideal $u^{\Delta}\overline{R}$.
Note that the ring $R$ is Gorenstein (being plane) and the symmetry of the equation (13) is the just the characterization of the Gorenstein condition of the ring $R$ (cf. [3]).
7 The Poincaré series of a divisorial valuation
Let $\nu=\nu_{\delta}$ be a divisorial valuation on ${\mathcal{O}}_{\mathbb{C}^{2},0}$. This means that it is defined by a component $E_{\delta}$ of the exceptional divisor of a modification of $(\mathbb{C}^{2},o)$. Let $\pi:(W,{\mathcal{D}})\to(\mathbb{C}^{2},0)$ be the minimal $\mathbb{K}$-resolution of the valuation $\nu$. The quotient of the dual graph of this resolution by the action of the Galois group $G$ looks like in figure 6.
Figure 6: The resolution graph $\check{\Gamma}$ for a divisorial valuation.
This graph essentially coincides with the minimal resolution graph of a $\mathbb{K}_{\delta}$-curvette at the component $E_{\delta}$ ($\mathbb{K}_{\delta}=\mathbb{K}_{s}$) with a possible tail from $\tau_{g}$ to $\delta$ added. Here, as above, $\sigma_{i}$, $i=0,1,\ldots,g$ are the dead ends of the graph $\check{\Gamma}$, $\tau_{i}$, $i=1,\ldots,g$ are the rupture points and $\rho_{j}$, $j=1,\ldots,s$ are the branching points of the $\mathbb{K}$-resolution.
Remark 11.
Pay attention that $\rho_{s}\neq\delta$. One may define $\rho_{s+1}$ as $\delta$. In this case the divisorial valuation $\nu_{\delta}$ coincides with the curve valuation $\nu_{C^{*}}$ for $C^{*}$ being a curvette at the component $E_{\delta}$ at a non-algebraic point of it. (Thus $\ell_{s+1}$ is assumed to be equal to $\infty$.)
We shall keep the notations ${\mathcal{E}}^{\{k_{\sigma}\}}_{\mathbb{K},0}$, $F^{\{k_{\sigma}\}}$, and $d^{\{k_{\sigma}\}}$ as before.
Theorem 3.
One has
$P_{\nu}(t)=P^{S}(t)\frac{1}{1-t^{M_{\delta}}}\cdot\prod_{i=1}^{s}\frac{1-t^{\ell_{i}M_{\rho_{i}}}}{1-t^{M_{\rho_{i}}}}\;.$ (14)
Proof..
Let $C$ be a $\mathbb{K}_{s}$-curvette at the component $E_{\delta}$. For the coefficient $a_{v}$ in the Poincaré series $P_{\nu}(t)=\sum a_{v}t^{v}$ one has $a_{v}:=\max d^{\{k_{\sigma}\}}$ where the maximun is taken over the collections $\{k_{\sigma}\}$ such that $\nu(\{k_{\sigma}\})=\sum k_{\sigma}M_{\sigma}=v$. Let $A^{\prime}$ be the set of collections $\{k_{\sigma}\}$, $\sigma\in\check{\Gamma}$ such that $k_{\delta}=0$ and let $P^{\prime}(t)=\sum a^{\prime}_{v}t^{v}$ with $a^{\prime}_{v}:=\max d^{\{k_{\sigma}\}}$ where the maximun is taken over the collections $\{k_{\sigma}\}\in A^{\prime}$ such that $\nu(\{k_{\sigma}\})=v$.
Lemma 9.
One has
$P^{\prime}_{\nu}(t)=P_{C}(t)=P^{S}(t)\cdot\prod_{j=1}^{s}\frac{1-t^{\ell_{j}M_{\rho_{j}}}}{1-t^{M_{\rho_{j}}}}\;.$
Proof..
The statement is a direct consequence of Remark 8. ∎
Lemma 10.
Let $v\in S^{\mathbb{K}}_{C}$ be of the form $v=\ell M_{\delta}+b$ with $b\in S^{\mathbb{K}}_{C}$ and the maximal possible $\ell$ in this representation. Then $a_{v}=\ell[\mathbb{K}_{s}:\mathbb{K}_{0}]+a^{\prime}_{b}$.
Proof..
(Cf. proof of Lemma 6.) Let us represent a collection $\{k_{\sigma}\}$ with $\nu(\{k_{\sigma}\})=v$ in the form $k_{\sigma}=k^{1}_{\sigma}+k^{2}_{\sigma}$, where $k^{1}_{\delta}=\ell^{\prime}$ ($\ell^{\prime}\leq\ell$), $k^{1}_{\sigma}=0$ for $\sigma\neq\delta$, $k^{2}_{\delta}=0$. A function $f\in{\mathcal{E}}_{\mathbb{K}^{2},0}^{\{k_{\sigma}\}}$ can be represented as $f=f_{1}f_{2}$ with $f_{r}\in{\mathcal{E}}_{\mathbb{K}^{2},0}^{\{k^{r}_{\sigma}\}}$, $r=1,2$. Let $h_{r}$, $r=1,2$, be fixed functions from ${\mathcal{E}}_{\mathbb{K}^{2},0}^{\{k^{r}_{\sigma}\}}$, $h=h_{1}h_{2}$. As in Lemma 6, we shall use a coordinate on $E_{\delta}$ such that the intersection point with the previous component from the geodesic to $\sigma_{0}$ is equal to $\infty$ and we shall take the function $h_{1}$ described there (that is, the $\ell^{\prime}$th power of the left hand side of the equation of the $G_{s}$-curvette at the point $w=1$ of the component $E_{\delta}$). One has $\Psi={\widetilde{f}}/{\widetilde{h}}=\Psi_{1}\Psi_{2}$, where $\Psi_{r}={\widetilde{f}_{r}}/{\widetilde{h}_{r}}$. The function $\Psi_{2}$ is constant on the component $E_{\delta}$. The maximal possible dimension (over $\mathbb{K}$) of the set of its values on this component is $a^{\prime}_{b^{\prime}}$ ($b^{\prime}=v-\ell^{\prime}M_{\delta}$). The function $\Psi_{1}$ is constant on the union of all the components $E_{\sigma}$ for $\sigma$ from the geodesic between $\sigma_{0}$ and $\delta$, excluding the last one. Its restriction to the component $E_{\delta}$ is equal to $c\prod_{p=1}^{\ell^{\prime}}(w-w_{p})/(w-1)^{\ell^{\prime}}$ with $c\in\mathbb{K}$, where $w_{p}\in\mathbb{K}_{s}$ are the intersection points of the strict transform of the curve $\{f_{1}=0\}$ with the component $E_{\delta}$. Therefore, up to the fixed multiple $(w-1)^{\ell^{\prime}}$, the restriction of $\Psi$ to the component $E_{\delta}$ is equal to the restriction of $\Psi_{2}$ multiplied by $c\prod_{p=1}^{\ell^{\prime}}(w-w_{p})$. The dimension (over $\mathbb{K}$) of the space of functions of the latter form is $\ell^{\prime}[\mathbb{K}_{s}:\mathbb{K}]$ (since $w_{p}$ are arbitrary elements of $\mathbb{K}_{s}$). Therefore the (maximal) dimension of the set of these functions on $E_{\delta}$ is $a^{\prime}_{b^{\prime}}+\ell^{\prime}[\mathbb{K}_{s}:\mathbb{K}_{0}]$. Thus it is maximal only for $\ell^{\prime}=\ell$ and is equal to $\ell[\mathbb{K}_{s}:\mathbb{K}_{0}]+a^{\prime}_{b}$. ∎
Just in the same way as Proposition 3 we get the following statement.
Proposition 4.
One has
$P_{\nu}(t)=P^{\prime}_{\nu}(t)\cdot\frac{1}{1-t^{M_{\delta}}}.$
Lemma 9 and Proposition 4 imply the statement of Theorem 3. ∎
Remark 12.
One can see that the classical Poincaré series of a curve or of a divisorial valuation determines the combinatorial type of the minimal $\mathbb{K}$-resolution of the valuation.